Mr qadri bought a piece of land for Rs 3,60,000 at Rs 1200 per sq. m. Find the length of the plot if the breadth is 3\4 of the length

Answer by mananth(16946)


Mr qadri bought a piece of land for Rs 3,60,000 at Rs 1200 per sq. m. Find the length of the plot if the breadth is 3\4 of the length
land for Rs 3,60,000 at Rs 1200 per sq. m
area of land = 360000/1200= 300 m^2
Let length l be x m
breadth will be 3x/4
Area = L * B
x * 3x/4 = 300
3x^2/4 =300
3x^2= 1200
x^2= 400
x= 20 m